# Spin-operator form factors of the critical Ising chain and their finite volume scaling limits

### v3

@@ Abstract
In this work, we provide a self-contained derivation of the spin-operator matrix elements in the fermionic basis, for the critical periodic Ising chain at a generic system length $N\in 2Z_{\geq 2}$. The approach relies on the near-Cauchy property of certain matrices formed by the Toeplitz symbols in the critical model, and leads to simpler product formulas for the dressing functions in terms of square root functions. These square root products allow fully dis-singularized integral representations. In the finite volume scaling limit, they further reduce to the Binet’s second integral for the gamma function logarithm and its Hermite’s generalization. As such, all the matrix elements in the scaling limit allow simple product formulas in terms of the gamma function at integer and half-integer arguments, and are rational numbers up to $\sqrt{2}$. They are exactly the spin-operator form factors of the Ising CFT in the fermionic basis, whose explicit forms are much less well known in comparison to the finite-volume form factors in the massive theory. We also fully determine the normalization factor of the spin-operator and show explicitly how the coefficient $G(\frac{1}{2})G(\frac{3}{2})$ appear through a ground state overlap. As a by-product, we obtain a Mellin-Barnes representation of the Glaisher constant with two $\zeta$ functions
$\int_{0<{\rm Re}(u)<1}\frac{du}{2\pi i}(2-2^{u})(2-2^{-u})\Gamma(u)\Gamma(-u)\zeta(u)\zeta(-u)=\ln\bigg(\pi^{\frac{1}{4}}G\left(\frac{1}{2}\right)G\left(\frac{3}{2}\right)\bigg)\ ,$ (1)
which allows to establish analytically the small mass asymptotics of a related crossover function.
@@ 1 The critical Ising chain, square-root products and main results
Similarly, in the literature, there are plenty discussions and derivations Lisovy (2002); Bugrij and Lisovyy (2003, 2004); Gehlen et al. (2008); Iorgov et al. (2011); Iorgov and Lisovyy (2011) for the finite volume form factors in the Ising model and the Ising chain. These works are very non-trivial, but most of them were assuming $T>T_{c}$ or $T<T_{c}$, and forgot to mention what happened at the critical point. For example, in the published version of Gehlen et al. (2008), the authors said explicitly after the Eq. (78) that the results are valid both when $g>1$ and $0\leq g<1$, but just omitted the critical point $g=1$. Very often, the formulas were written with a refactor such as $\xi=(g^{2}-1)^{\frac{1}{4}}$ in front, a factor that vanishes at the critical point $g=1$, like the expressions given to the leg functions in the massive theory, despite that in the proper way of taking the $g\rightarrow 1^{+}$ limit, the results are actually finite and as we will show, lead exactly to the form factors at the critical point.
Meanwhile, in the literature, there were also not too many discussions on the spin-operator form factors in the Ising-CFT, despite that they are clearly universal, beautiful and unique quantities: the massless fermion basis of the Ising CFT is perhaps among the few integrable massless scattering basis Bazhanov et al. (1996, 1997) of 2D CFTs, that allow explicit expressions for local operator’s matrix elements between them. On the other hand, in comparison to the Virasoro basis, it could be the massless scattering basis, that respect better the scale-full deformations of CFTs either to the integrable lattice models in the UV Boos et al. (2010), or to the massive QFTs in the IR. It is possible, that for certain non-inclusive high-energy limits that are more sensitive to the structures of the Hilbert space, one needs not just the standard CFT perturbation theory for point-like objects, but also the correct massless basis, to write out the correct high-energy asymptotics. These questions deserve more investigations, and were among the motivations of this work.
Given all above, the author believes that it is reasonable and necessary to provide a self-contained derivation for the spin-operator form factors in the 2D periodic Ising chain, directly at the critical point. Below we will provide such a derivation, and show how in the finite volume scaling limit, the results converge to the Ising CFT versions. We believe that the products, integral representations and the large $N$ asymptotics in this work are simple yet beautiful. We also believe that the explicit formulas provided by this work could be of practical usages (such as the TFCSA applications Yurov and Zamolodchikov (1991)).
@@ 1.1 The square root function and roots of $\pm 1$
They are the exponentials of the periodic and anti-periodic momenta. Since $N$ is even, $1\in S_{+}$, and points in $S_{+}$ except 1,11,-1 can be grouped to $\frac{N-2}{2}$ pairs symmetric with respect to the real axis. Similarly, all points in $S_{-}$ are grouped to $\frac{N}{2}$ such pairs. Using these properties, one has first of two of the product formulas to be used latter
@@
Latter we will use these formula to simply the leg functions at $\beta=1$.
@@ 1.2 The critical Ising chain
$H=\frac{i}{2}\sum_{i=0}^{2N-2}\Gamma_{i}\Gamma_{i+1}-\frac{i}{2}\Gamma_{2N-1}\Gamma_{0}Z_{2}\ ,\ Z_{2}=\prod_{i=0}^{N-1}(-i)\Gamma_{2i}\Gamma_{2i+1}\ .$ (13)
The operator $Z_{2}$ is just the standard $Z_{2}$ charge and commutes with $H$ and other even combinations of fermion operators. As such, in the charge odd and even sectors, one has two different quadratic Hamiltonians, but in terms of the same set of fermion operators. In each of the sectors, one can diagonalize the Hamiltonian by finding the proper, orthogonal transformations that diagonalize the skew-symmetric matrices appearing in $H$.
To further set up our conventions for the eigensystem, here we review how the diagonalization process works in the $Z_{2}=-1$ (R) sector. The $Z_{2}=1$ (NS) sector can be handled similarly and is simpler. We will mainly follow the spirit of Kaufman (1949), but the computations are simpler in the Ising chain than the Ising model. For $Z_{2}=-1$, one has the following rotation matrix
@@
$H=\frac{i}{4}\sum_{i,j}\Gamma_{i}\Gamma_{j}{\cal E}_{ij}\ .$ (15)
@@
The factor $\frac{1}{\sqrt{N}}$ is because the norm of the vector $\lambda(\beta)$ is $\sqrt{2N}$. From the rotation matrix, one can form the rotated Dirac matrices
@@
that block diagonalize the Hamiltonian
$H=\frac{i}{2}\sum_{\beta\in S_{+}}\epsilon(\beta)\gamma_{\beta}\gamma_{\beta+\frac{1}{2}}\ .$ (27)
At this point, one only knows that $\Omega^{T}\Omega=1$, but to further determine the spectrum, one needs to know the sign of $\det\Omega$.
@@
$\varphi(\beta,h)=i\left(\frac{1-h\beta}{1-h\bar{\beta}}\right)^{\frac{1}{2}}\ ,\ 0\leq h\leq 1\ ,$ (28)
@@
justifying the choice made above. The continuity implies that $\det\Omega(1^{-})=\det\Omega(0)$, since the determinant is a continuous function of $h$ which could take only two values $\pm 1$. It is not hard to compute that $\det\Omega(0)=1$, since at $h=0$ the Hamiltonian is non-interacting, and $\Omega(0)$ is a $2N\times 2N$ rotation underlying an $N\times N$ unitary transformation. As such, the rotation matrix $\Omega=\Omega(1^{-})$ is already proper. Notice if one choose $\varphi(1)=-i$, then the $\det\Omega$ would be 1-1, since this choice only changes the sign of $\gamma_{0}$.
Knowing the sign of $\det\Omega$, one can further introduce the creation and annihilation operators
@@
in terms which the diagonalized Hamiltonian and the charge operator read
$H=\sum_{\beta\in S_{+}}\epsilon(\beta)\left(c_{\beta}^{\dagger}c_{\beta}-\frac{1}{2}\right)\ ,$ (31)
$Z_{2}=\det\Omega\prod_{\beta\in S_{+}}(-i\gamma_{\beta}\gamma_{\beta+\frac{1}{2}})=\prod_{\beta\in S_{+}}(1-2c_{\beta}^{\dagger}c_{\beta})\ .$ (32)
Since $\Omega$ is non-degenerate, one can always find a normalized state $|\Omega_{R}\rangle\neq 0$ annihilated by all the $c_{\beta}$s. This state must be charge even, due to the Eq. (32) for the $Z_{2}$ charge. Since the representation space is $2^{N}$ dimensional, the Fock space generated by acting the $c_{\beta}^{\dagger}$s on $|\Omega_{R}\rangle$ is the full space. Among them, only those with an odd number of $c_{\beta}^{\dagger}$s are energy eigenstates, due to the $Z_{2}=-1$ condition. The diagonalization process in the NS sector is similar. All the expressions in the R sector remain the same form, except $\beta\in S_{+}$ should be replaced by $z\in S_{-}$. There is no $z=1$ issue in NS and $\varphi(z)=\sqrt{z}$ for all $z\in S_{-}$.
To summarize, in terms of the following Toeplitz symbol [^2] The choice is to be consistent with the symbol $C(z,h)=\left(\frac{1-h\bar{z}}{1-hz}\right)^{\frac{1}{2}}$ for the Toeplitz determinant of the $\sigma\sigma$ two point correlator in the ordered region.
@@
one has the mode decompositions in terms of $\beta\in S_{+}$
@@
and similarly in term of $z\in S_{-}$
@@
The Hamiltonian in the NS sector is diagonalized by $c_{z},c_{z}^{\dagger}$, and in the R-sector is diagonalized by $c_{\beta},c_{\beta}^{\dagger}$. The state $|\Omega_{R}\rangle$ annihilated by all the $c_{\beta}$s is charge even. Meanwhile, the ground state $|\Omega_{NS}\rangle$ annihilated by all the $c_{z}$s is also charge even and is the absolute ground state of the whole system, whose energy will be normalized to 0. A complete set of energy eigenstates of the system and their energies are therefore
Notice that $c_{1}^{\dagger}|\Omega_{R}\rangle$ is the physical ground state in the charge odd sector, and the $c_{1}^{\dagger}$ excitation is called the zero mode. In the diagonalized Hamiltonian, $c_{1}$ and $c_{1}^{\dagger}$ disappear, but the energy difference $\Delta_{E}(N)$ between the R and NS ground states is non-zero and approaches $\frac{\pi}{4N}$ as $N\rightarrow\infty$. To show this, one simply notice that
The sums are geometric and can be performed in elementary manner. However, to illustrate the method we will use to perform such sums for more complicated functions, it is also possible to use the weighting functions $\frac{1}{2\pi iz}\frac{-N}{z^{N}+1}$ and $\frac{1}{2\pi i\beta}\frac{N}{\beta^{N}-1}$ to express the summations as contour integrals, and then deform the contours to the branch cut along the positive real axis. At the end, after partial integration one has the following integral representation
This $\frac{\pi}{4N}$ is exactly due to the scaling dimension of the spin operator and allows the identification of the state $c_{1}^{\dagger}|\Omega_{R}\rangle$ as the primary state $|\hat{\sigma}\rangle$ of the spin operator in the CFT language.
We should mention that although $|\Omega_{R}\rangle$ is not an eigenstate of the $H$ in Eq. (8), it is the absolute ground state of another “twisted” Hamiltonian which relates to $H$ by a sign flip in the $\sigma^{x}_{N-1}\sigma^{x}_{0}$ term
@@
Indeed, the fermion quadratic form for $H^{\prime}$ simply differs from Eq. (13) by a sign flip in the $Z_{2}$ dependent term. As such, the spectrum for $H^{\prime}$ can still be expressed in terms of the same sets of $c_{z}^{\dagger},c_{\beta}^{\dagger}$s as $H$, but with even numbers of excitations on top of $|\Omega_{R}\rangle$, and odd numbers of excitations on top of $|\Omega_{NS}\rangle$. As such, it is reasonable to call $|\Omega_{R}\rangle$ also as a ground state, and the quantity $\langle\Omega_{R}|\Omega_{NS}\rangle$ as a ground state overlap. This observation also helps to establish the property of the eigensystem Eq. (38) and Eq. (39) under the spatial translation. For this, we introduce following translation operations
@@
$T_{\pm}\Gamma_{i}=\Gamma_{i+2}T_{\mp},\ 0\leq i\leq 2N-3\ ,\ T_{\pm}\Gamma_{2N-2}=\pm\Gamma_{0}T_{\mp}Z_{2}\ ,T_{\pm}\Gamma_{2N-1}=\pm\Gamma_{1}T_{\mp}Z_{2}\ .$ (46)
@@
$T_{\pm}c_{z}^{\dagger}\frac{1\mp Z_{2}}{2}=zc_{z}^{\dagger}T_{\mp}\frac{1\mp Z_{2}}{2}\ ,\ T_{\pm}c_{\beta}^{\dagger}\frac{1\pm Z_{2}}{2}=\beta c_{\beta}^{\dagger}T_{\mp}\frac{1\pm Z_{2}}{2}\ .$ (47)
@@
As such, when $\langle\Omega_{R}|\Omega_{NS}\rangle\neq 0$, due to the reality property, the phase difference $e^{-i(\theta_{+}-\theta_{-})}$ must be $\pm 1$ and is an absolute quantity independent of the state normalization. Latter we will show explicitly using the overlapping matrix $v_{z\beta}$ define in Eq. (131) that $e^{i(\theta_{+}-\theta_{-})}=1$. Equivalently, the left continuity at $h=1$ and the absence of discontinuities when $h\in[0,1)$ allows $e^{i(\theta_{+}-\theta_{-})}$ to be computed by the $h=0$ version, which is clearly $1$. Indeed, when $h=0$, $H=H^{\prime}$, and the absolute ground state is a trivial “all-down state”.

### v4

@@ Abstract
In this work, we provide a self-contained derivation of the spin-operator matrix elements in the fermionic basis, for the critical periodic Ising chain at a generic system length $N\in 2Z_{\geq 2}$. The approach relies on the near-Cauchy property of certain matrices formed by the Toeplitz symbols in the critical model, and leads to a few square-root products for the leg functions. The square root products allow simple integral representations, that further reduce to the Binet’s second integral and its generalization by Hermite, in the finite volume scaling limit. This leads to product formulas for the spin operator matrix elements in the scaling limit, providing explicit expressions for the spin-operator form factors of the Ising CFT in the fermionic basis, that were computed iteratively in Yurov and Zamolodchikov (1991). They are all rational numbers up to $\sqrt{2}$. We also determine the normalization factor of the spin-operator and show explicitly how the coefficient $G(\frac{1}{2})G(\frac{3}{2})$ appear through a ground state overlap. Moreover, by expanding the spin-spin two point correlator in the fermionic basis, we observed a Fredholm determinant identity
$\det(1+{\cal K}_{ij}(w))_{1\leq i,j\leq\infty}=(1-w)^{-\frac{1}{8}},\ \ {\cal K}_{ij}(w)=\frac{1}{\pi}\frac{i}{i+j}\frac{\Gamma(i+\frac{1}{2})\Gamma(j+\frac{1}{2})}{\Gamma(i+1)\Gamma(j+1)}w^{j}\ ,$ (1)
which allows to show the convergence of the rescaled two-point correlator to the CFT version on a cylinder.
@@ 1 The critical Ising chain, square-root products and main results
Similarly, in the literature, there are plenty discussions and derivations Bugrij (2001); Lisovy (2002); Bugrij and Lisovyy (2003, 2004); Gehlen et al. (2008); Iorgov et al. (2011); Iorgov and Lisovyy (2011) for the finite volume form factors in the Ising model and the Ising chain. These works are very non-trivial, but most of them were assuming $T>T_{c}$ or $T<T_{c}$, and forgot to mention what happened at the critical point. For example, in the published version of Gehlen et al. (2008), the authors said explicitly after the Eq. (78) that the results are valid both when $g>1$ and $0\leq g<1$, but just omitted the critical point $g=1$. Very often, the formulas were written with a refactor such as $\xi=(g^{2}-1)^{\frac{1}{4}}$ in front, a factor that vanishes at the critical point $g=1$, like the expressions given to the leg functions in the massive theory, despite that in the proper way of taking the $g\rightarrow 1^{+}$ limit, the results are actually finite and as we will show, lead exactly to the form factors at the critical point.
Meanwhile, in the literature, there were also not too many discussions on the spin-operator form factors in the Ising-CFT, despite that they are clearly universal and unique quantities: the massless fermion basis of the Ising CFT is perhaps among the few integrable massless scattering basis Bazhanov et al. (1996, 1997) of 2D CFTs, that allow explicit expressions for local operator’s matrix elements between them. On the other hand, in comparison to the Virasoro basis, it could be the massless scattering basis, that respect better the scale-full deformations of CFTs either to the integrable lattice models in the UV Boos et al. (2010), or to the massive QFTs in the IR. It is also possible, that for certain non-inclusive high-energy limits that are more sensitive to the structures of the Hilbert space, one needs not just the standard CFT perturbation theory for point-like objects, but also the correct massless basis, to write out the correct high-energy asymptotics. These questions deserve more investigations. Furthermore, to the author’s taste, CFTs are beautiful not just for themselves, but also for the explicit asymptotics telling how they are approached. It is always interesting to establish such asymptotics from the first principle for various quantities, whenever possible.
Given all above, the author believes that it is not unreasonable to provide a self-contained derivation for the spin-operator form factors in the 2D periodic Ising chain, directly at the critical point. Below we will provide such a derivation, and show how in the finite volume scaling limit, the results converge to the Ising CFT versions. We believe that many of the products, integral representations, large $N$ asymptotics and the matrix elements in this work are novel, beautiful and simple. We also believe that the explicit formulas provided by this work could be of practical usages (such as the TFCSA applications Yurov and Zamolodchikov (1991)).
@@ 1.1 The square root function and roots of $\pm 1$
They are the exponentials of the periodic and anti-periodic momenta. Throughout this paper, the small letter $\beta$ in formulas will be reserved for $S_{+}$, and the small letter $z$ reserved for $S_{-}$. Since $N$ is even, $1\in S_{+}$, and points in $S_{+}$ except 1,11,-1 can be grouped to $\frac{N-2}{2}$ pairs symmetric with respect to the real axis. Similarly, all points in $S_{-}$ are grouped to $\frac{N}{2}$ such pairs. Using these properties, one has first of two of the product formulas to be used latter
@@
Latter we will use these formula to simplify the leg functions at $\beta=1$.
@@ 1.2 The critical Ising chain
$H=\frac{i}{2}\sum_{i=0}^{2N-2}\Gamma_{i}\Gamma_{i+1}-\frac{i}{2}\Gamma_{2N-1}\Gamma_{0}Z\ ,\ Z=\prod_{i=0}^{N-1}(-i)\Gamma_{2i}\Gamma_{2i+1}\ .$ (13)
The operator $Z$ is just the standard $Z_{2}$ charge and commutes with $H$ and other even combinations of fermion operators. As such, in the charge odd and even sectors, one has two different quadratic Hamiltonians, but in terms of the same set of fermion operators. In each of the sectors, one can diagonalize the Hamiltonian by finding the proper, orthogonal transformations that diagonalize the skew-symmetric matrices appearing in $H$.
To further set up our conventions for the eigensystem, especially, to convince the readers that our sign choices at the $\beta=1$ point are correct, here we review how the diagonalization process works in the Z=1Z=-1 (R) sector. The $Z=1$ (NS) sector can be handled similarly and is free from the $\beta=1$ issues. We will mainly follow the spirit of Kaufman (1949), but the computations are simpler in the Ising chain than the Ising model. For Z=1Z=-1, one has the following rotation matrix
@@
$H(1-Z)=\frac{i}{4}\sum_{i,j}\Gamma_{i}\Gamma_{j}{\cal E}_{ij}(1-Z)\ .$ (15)
@@
The factor $\frac{1}{\sqrt{N}}$ is because the norm of the unnormalized vector $\lambda(\beta)$ in Eq. (18) is $\sqrt{2N}$. From the rotation matrix, one can form the rotated Dirac matrices
@@
that block diagonalize the Hamiltonian under the 1Z1-Z projection
$H\frac{1-Z}{2}=\frac{i}{2}\sum_{\beta\in S_{+}}\epsilon(\beta)\gamma_{\beta}\gamma_{\beta+\frac{1}{2}}\frac{1-Z}{2}\ .$ (27)
At this point, one only knows that $\Omega^{T}\Omega=1$, but to further determine the spectrum, one needs to know the sign of $\det\Omega$ in order to express the charge operator $Z$ in terms of the $\gamma_{\beta},\gamma_{\beta+\frac{1}{2}}$s.
@@
$\varphi(\beta,h)=i\left(\frac{1-h\beta}{1-h\bar{\beta}}\right)^{\frac{1}{2}}\ ,\ 0\leq h<1\ ,$ (28)
@@
justifying the choice made above. The continuity implies that $\det\Omega(1^{-})=\det\Omega(0)$, since the determinant is a continuous function of $h$ which could take only two values $\pm 1$. It is not hard to compute that $\det\Omega(0)=1$, since at $h=0$ the Hamiltonian is non-interacting, and $\Omega(0)$ is a $2N\times 2N$ rotation underlying an $N\times N$ unitary transformation. As such, the rotation matrix $\Omega=\Omega(1^{-})$ is already proper. Notice if one choose $\varphi(1)=-i$, then the $\det\Omega$ would be 1-1, since this choice only changes the sign of $\gamma_{0}$. Knowing the sign of $\det\Omega$, one can further introduce the creation and annihilation operators
@@
in terms which the Hamiltonian under the 1Z1-Z projection and the charge operator read
$H\frac{1-Z}{2}=\sum_{\beta\in S_{+}}\epsilon(\beta)\left(c_{\beta}^{\dagger}c_{\beta}-\frac{1}{2}\right)\frac{1-Z}{2}\ ,$ (31)
$Z=\det\Omega\prod_{\beta\in S_{+}}(-i\gamma_{\beta}\gamma_{\beta+\frac{1}{2}})=\prod_{\beta\in S_{+}}(1-2c_{\beta}^{\dagger}c_{\beta})\ .$ (32)
Since $\Omega$ is non-degenerate, one can always find a normalized state $|\Omega_{R}\rangle\neq 0$ annihilated by all the $c_{\beta}$s. This state must be charge even, due to the Eq. (32) for the $Z_{2}$ charge. Since the representation space is $2^{N}$ dimensional, the Fock space generated by acting the $c_{\beta}^{\dagger}$s on $|\Omega_{R}\rangle$ is the full space. Among them, only those with an odd number of $c_{\beta}^{\dagger}$s are energy eigenstates, due to the 1Z1-Z projection. The diagonalization process in the NS sector is similar. All the expressions in the R sector remain the same form, except $\beta\in S_{+}$ should be replaced by $z\in S_{-}$. There is no $z=1$ issue in NS and $\varphi(z)=\sqrt{z}$ for all $z\in S_{-}$.
Combining the two sectors, the full Hamiltonian reads
$H=\sum_{\beta\in S_{+}}\epsilon(\beta)\left(c_{\beta}^{\dagger}c_{\beta}-\frac{1}{2}\right)\frac{1-Z}{2}+\sum_{z\in S_{-}}\epsilon(z)\left(c_{z}^{\dagger}c_{z}-\frac{1}{2}\right)\frac{1+Z}{2}\ ,$ (33)
$Z=\prod_{\beta\in S_{+}}(1-2c_{\beta}^{\dagger}c_{\beta})=\prod_{z\in S_{-}}(1-2c_{z}^{\dagger}c_{z})\ .$ (34)
The state $|\Omega_{NS}\rangle$ annihilated by all the $c_{z}$s is the ground state in the charge even sector, sector, and the $c_{1}^{\dagger}|\Omega_{R}\rangle$ is the ground state in the charge odd sector. The $c_{1}^{\dagger}$ excitation will be called the zero mode. On the other hand, although $\epsilon(1)=0$, the energy difference $\Delta_{E}(N)$ between the R and NS ground states is non-zero and approaches $\frac{\pi}{4N}$ as $N\rightarrow\infty$. To show this, notice that
The sums are geometric and can be performed in elementary manner. It is also possible to use the weighting functions $\frac{1}{2\pi iz}\frac{-N}{z^{N}+1}$ and $\frac{1}{2\pi i\beta}\frac{N}{\beta^{N}-1}$ to express the summations as contour integrals, and then deform the contours to the branch cuts along the positive real axis. At the end, one has
As such, the state $|\Omega_{NS}\rangle$ is the absolute ground state of the system, whose energy will be normalized to 0. On the other hand, the energy gap $\frac{\pi}{4N}$ is exactly due to the scaling dimension of the spin operator and allows the identification of the state $c_{1}^{\dagger}|\Omega_{R}\rangle$ as the primary state $|\hat{\sigma}\rangle$ of the spin operator in the CFT language. On top of the ground states, a complete set of energy eigenstates of the system and their energies are given by
They are exactly the fermionic basis, in terms which we will compute the matrix elements of the spin operator.
Equally important as the spectrum for the form factor computations are the mode decompositions of the $\Gamma_{i}$s. In terms of the following Toeplitz symbol [^2] The choice is to be consistent with the symbol $C(z,h)=\left(\frac{1-h\bar{z}}{1-hz}\right)^{\frac{1}{2}}$ for the Toeplitz determinant of the $\sigma\sigma$ two point correlator in the ordered region.
@@
the mode decompositions in terms of $c_{\beta},c_{\beta}^{\dagger}$ follow by inverting the Eq. (25), Eq. (26) using $\Omega\Omega^{T}=1$
@@
Similarly, the same set of $\Gamma_{i}$s can also be expressed in term of $c_{z},c_{z}^{\dagger}$
@@
To obtain the form factors of the charge odd operators such as the spin operator, one needs to compute the matrix elements between the two Fock basis defined by two sets of creation-annihilation operators Abraham (1978a, b). In the continuum limit, this is not a well posed problem. But in the finite $2^{N}$ dimensional Hilbert space, since both of the $c_{z},c_{z}^{\dagger}$ and $c_{\beta},c_{\beta}^{\dagger}$ are ultimately defined linearly by the same sets of $\Gamma_{i}$s, this problem is well posed and can be reduced to matrix inversion and multiplication problems in the $N\times N$ one-particle Hilbert space. We will show in Sec. 2, that the matrices to be inverted and multiplied at the critical point are rather special. As a result, elementary derivations of the form factors exist.
Before moving to the form factors, let’s stay in this introductory subsection a little bit longer and address the property of the eigensystem Eq. (37) and Eq. (38) under the spatial translation. The first observation to make is, although $|\Omega_{R}\rangle$ is not an eigenstate of the $H$ in Eq. (8), it is the absolute ground state of another “twisted” Hamiltonian which relates to $H$ by a sign flip in the $\sigma^{x}_{N-1}\sigma^{x}_{0}$ term
@@
Indeed, the fermion quadratic form for $H^{\prime}$ simply differs from Eq. (13) by a sign flip in the $Z_{2}$ dependent term. As such, the spectrum for $H^{\prime}$ can still be expressed in terms of the same sets of $c_{z}^{\dagger},c_{\beta}^{\dagger}$s as $H$, but with even numbers of excitations on top of $|\Omega_{R}\rangle$, and odd numbers of excitations on top of $|\Omega_{NS}\rangle$. We then introduce the following translation operations
@@
$T_{\pm}\Gamma_{i}=\Gamma_{i+2}T_{\mp},\ 0\leq i\leq 2N-3\ ,\ T_{\pm}\Gamma_{2N-2}=\pm\Gamma_{0}T_{\mp}Z\ ,T_{\pm}\Gamma_{2N-1}=\pm\Gamma_{1}T_{\mp}Z\ .$ (48)
@@
$T_{\pm}c_{z}^{\dagger}\frac{1\mp Z}{2}=zc_{z}^{\dagger}T_{\mp}\frac{1\mp Z}{2}\ ,\ T_{\pm}c_{\beta}^{\dagger}\frac{1\pm Z}{2}=\beta c_{\beta}^{\dagger}T_{\mp}\frac{1\pm Z}{2}\ .$ (49)
@@
As such, when $\langle\Omega_{R}|\Omega_{NS}\rangle\neq 0$, due to the reality property, the phase difference $e^{-i(\theta_{+}-\theta_{-})}$ must be $\pm 1$ and is an absolute quantity independent of the state normalization. Latter we will show explicitly using the overlapping matrix $v_{z\beta}$ define in Eq. (133) that $e^{i(\theta_{+}-\theta_{-})}=1$. Equivalently, the left continuity at $h=1$ and the absence of discontinuities when $h\in[0,1)$ allow the $e^{i(\theta_{+}-\theta_{-})}$ to be computed by the $h=0$ version, which is clearly $1$. Indeed, when $h=0$, $H=H^{\prime}$, and the absolute ground state is a trivial “all-down state”.